( 1 /t + 1/ t ^2 − y /t ^2 + y ^2 ) d t + ( y e ^y + t /t ^2 + y^ 2 ) d y = 0


1
Expert's answer
2021-07-18T08:58:39-0400

"( {1 \over t} + {1\over t ^2} − {y \over t ^2 + y ^2 }) d t + ( y e ^y +{ t \over t ^2 + y^ 2 }) d y = 0"


"M(t,y)=" "( {1 \over t} + {1\over t ^2} − {y \over t ^2 + y ^2 })"


"\implies { \partial M(t,y)\over \partial y}=M_y=-{1(t^2+y^2)-2y(y) \over (t^2+y^2)^2}=-{t^2+y^2-2y^2\over (t^2+y^2)^2}=-{t^2-y^2\over (t^2+y^2)^2}={y^2-t^2\over (t^2+y^2)^2}"


"N(t,y)=( y e ^y +{ t \over t ^2 + y^ 2 })"


"\implies { \partial N(t,y)\over \partial t}=N_t={y^2-t^2\over (t^2+y^2)^2}"


The differential equation is exact since "M_y=N_t"


There exist a function "\mu(t,y)" such that


"{\partial \mu (t,y)\over \partial t}=M(t,y)=( {1 \over t} + {1\over t ^2} − {y \over t ^2 + y ^2 })"


"{\partial \mu (t,y)\over \partial y}=N(t,y)=( y e ^y +{ t \over t ^2 + y^ 2 })"


Let "\mu (t,y)=\int^yN(t,y)dy" "+" "g(t)"


"=\int^y( y e ^y +{ t \over t ^2 + y^ 2 }) dy" "+" "g(t)"


"=ye^y-e^y+tan^{-1}({y\over t})+g(t)"


"\implies {\partial \mu (t,y)\over \partial t}=-{y\over t^2+y^2}+{\partial g(t)\over \partial t}={1 \over t} + {1\over t ^2} − {y \over t ^2 + y ^2 }"


"\implies {\partial g(t)\over \partial t}={1 \over t} + {1\over t ^2}"


"\implies \int dg(t)=\int( {1 \over t} + {1\over t ^2})dt"


"\implies g(t)=ln(t)-{1\over t}"


"\implies \mu(t,y)=ye^y-e^y+tan^{-1}({y\over t})+ln(t)-{1\over t}"


"\therefore" The general solution is thus


"ye^y-e^y+tan^{-1}({y\over t})+ln(t)-{1\over t}=c"